# Elastic Shape Analysis of Movement Data

### v2

@@ 3.1.1 Implementation and penalty parameter selection
To address the lack of a common underlying structure, we registered the full dataset using the penalized elastic shape analysis procedure. By adjusting the elasticity parameter $\lambda$, we aligned the full set of GRF curves without distorting the shape of the atypical curves. To determine an appropriate $\lambda$, we computed warping functions iteratively over a grid of candidate values and visually examined the alignment results, focusing particularly on atypical curves. We refined the grid in regions where $\lambda$ values produced reasonable results, selecting values that avoided excessive smoothing of features in the atypical curves while still capturing meaningful phase variation. Over-alignment led to sharp corners forming a staircase-like pattern in the warping functions, indicating drastic stretching and compression of curves with differing underlying structures, such as unimodal curves.
@@
Based on this evaluation, we selected $\lambda=2$ as a reasonable balance between alignment and over-alignment. To illustrate this trade-off, Figure 3 compares results for three $\lambda$ values: the fully elastic case $\lambda=0$ in the top row, our selected $\lambda=2$ in the middle row, and a less elastic case $\lambda=4$ in the bottom row. Atypical cases are colored by speed, while other curves are shown in gray, as in Figure 2. The leftmost panels display the warping functions for each $\lambda$, where $\lambda=0$ results in sharp staircase-like patterns indicative of over-alignment, while $\lambda=4$ yields warping functions tightly clustered around the identity, suggesting insufficient alignment. The second through fourth columns show the aligned curves. Results for additional values of $\lambda$ can be found in Figure 11 of the supplementary material.
@@
While our approach treated alignment as agnostic to a specific performance criterion, a more formal selection method could be employed depending on the analytical objectives. For instance, if alignment quality is evaluated based on a predictive or inferential task, $\lambda$ could be optimized via cross-validation to maximize a relevant performance metric. However, in finite samples, cross-validation may yield a noisy selection due to its slow convergence to optimal results (see Hall and Marron (1987) in the context of bandwidth selection for kernel density estimation). Alternatively, if the goal is to estimate an underlying common signal, Kim et al. (2023) provides a scale-space approach for estimating both the shape of the unknown signal and the signal itself. Defining a notion of optimality depends on the specific context of the analysis, and the choice of $\lambda$ for different analytical objectives remains an open area of research.
@@
Figure 3: Alignment results for three elasticity parameter values: fully elastic ($\lambda=0$), the selected $\lambda=2$, and a more rigid case ($\lambda=4$). Atypical cases are colored by speed, while other curves are shown in gray. The leftmost panels display the warping functions, and the second through fourth columns show the aligned curves. The staircasing effect in the fully elastic case ($\lambda=0$) indicates over-alignment, while $\lambda=4$ results in warping functions tightly clustered around the identity, suggesting insufficient alignment. The selected $\lambda=2$ balances these effects.
@@
Next, we obtained amplitude objects by applying the (common) set of warping functions to each set of original GRF curves, which gives an intuitive representation of each component’s amplitude. The amplitude objects obtained for different values of $\lambda$ are shown in the second through fourth columns of Figure 3.
@@ 3.2 Modes of variation
In the OODA terminology, a collection of members of the object space that summarize one component of variation and is in some sense one-dimensional is called a mode of variation. For example, in the vector matrix case, a mode of variation is a rank-one matrix. We can obtain modes of variation through Principal Component Analysis (PCA), where each object is considered as a point in high dimensional space (column vector). For an introduction to PCA, see Jolliffe (2002). For each direction of the ground reaction force (vGRF, apGRF, mlGRF), PCA was performed on the set of 2686 Fisher-Rao aligned curves, each corresponding to an individual gait observation, yielding the amplitude modes (Tucker et al., 2013). The sets of input curves are shown in the second, third, and fourth columns of the middle row of Figure 3.
@@
Figure 4 shows the modes of variation of the amplitude objects of the vGRF, where the curves are colored according to walking speed. The first mode of variation, shown in the first panel of the second row of that figure, is associated with walking speed and reflects the contrast in peak heights and valley depths. Faster walkers (indicated in red in the rainbow color scheme) generally exhibit higher peaks and lower valleys, while slower walkers (purple in the rainbow color scheme) have lower peaks and a shallower valley. The middle column of that figure displays the largest (dashed curve) and smallest (dotted curve) PC projections added back to the mean curve, which is shown as a solid black curve in each of the middle panels. The middle panel of the first mode shows that the slowest walkers (dotted curve) exhibit a vertical amplitude force that appears unimodal and does not exceed body weight (1 on the vertical axis), indicating that these walkers do not fully transfer their weight to the striking limb. This type of gait can be thought of as “shuffling.” The second and third modes of variation are about the second and first peak, respectively. The middle panel in the third row shows the largest and smallest PC2 projections added back to the mean curve, distinguished with a dashed (largest) and dotted (smallest) line type. These extremes show that variation in this mode is mostly in the height of the second peak. In the panel below, the extremes of the third mode indicate phase variation in the first peak that is unique to the vertical component. The fourth mode of variation reflects the overall magnitude, particularly in the mid-stance phase. The second, third, and fourth PC projection extreme curves all suggest that some curves have a small third bump before the first peak. In gait analysis, this pattern is known as the heelstrike transient (HST), a rapid and transient rise in the vGRF immediately after ground contact. As discussed in Blackburn et al. (2016) and references therein, the presence and characterization (e.g. magnitude) of HST indicate impulsive loading, which influences cartilage degradation and symptoms of OA. However, Blackburn et al. (2016) also noted that methods for identifying HST can be unreliable. The amplitude modes of variation we identified offer a potentially viable method for reliably identifying and understanding the HST.
@@
We applied the PNS-based functional PCA methodology to the set of (common) warping functions to obtain phase modes of variation. We found that the great sphere decomposition from PNS yielded the most interpretable phase modes of variation because of weak interpretably of small sphere variation. Figure 5 depicts an intuitively useful notion of phase variation represented by warpings of the Karcher mean of the vGRF curves. The warping functions used to create these visualizations were generated by taking the inverse of the phase PNS projections added to the 45-degree line (identity warp). In each panel of the figure, the curves are colored based on the PNS scores for the corresponding mode, with cyan indicating the lowest scores and magenta indicating the highest. It is important to note that the curves are plotted in the order of the corresponding score, as over-plotting is an issue. The first mode (first panel) shows an overall shift in timing, with most apparent differences in the timing of the first peak (maximum heel-strike force) and valley. The second mode in the next panel appears to explain variability in the closeness of the peaks - the cyan curves are the curves with peaks closer together and the magenta curves have peaks farther apart. The third mode represents an overall phase shift (left vs. right) and seems to suggest that curves having a small third bump before the first peak correspond with an earlier timing (cyan curves), especially an earlier second peak. The fourth mode appears to highlight variability in the timing of the second peak, independent of the rest of the curve.
@@ 4 Comparison to conventional discrete summaries
We investigated the added value of analyzing patterns across the entire movement curve, as opposed to conventional discrete summaries of GRF curves. To assess this, we compared the predictive performance of predictors derived from our shape-based analysis of the full movement curves with conventional discrete summaries in predicting disease severity.
Osteoarthritis (OA) severity is a critical outcome in disease treatment studies and is commonly evaluated using joint space width (JSW) measurements, as well as the Kellgren-Lawrence grade (KLG), a radiographic classification system, and patient-reported pain and function scores from the Western Ontario and McMaster Universities Arthritis Index (WOMAC) (Alexandersen et al., 2014). JSW, measured in millimeters on radiographic images, is particularly important for diagnosing and monitoring knee OA.
In the IDEA study, JSW was available for only a subset of patients. Therefore, we analyze it separately in this section, while analyzing other OA measures with larger sample sizes in the following section. While JSW narrowing is a key marker of disease progression, prior studies suggest that knees at similar stages of OA tend to progress at comparable rates (Benichou et al., 2010). Consequently, baseline JSW may provide a more meaningful indicator of disease severity than JSW change over time. Furthermore, because the IDEA study is a clinical trial of interventions for patients already having an advanced stage of knee OA, focusing on baseline measures is the most appropriate approach for this dataset. Given these considerations, we used baseline JSW as a dependent variable to compare full-curve predictors with conventional discrete summary predictors.
Baseline JSW measurements were missing for 129 out of the 454 IDEA participants. The missing values for JSW are shown in the last column of the heatmap in Figure 15 in the supplementary material, with red lines indicating missing values. The rest of that figure is discussed in Section 4.2. The reasons for the missing values were not reported in the data, but a comparison of clinical data for participants with and without baseline JSW measurements showed no apparent sampling bias. This left a total of 325 participants with baseline JSW data for our analysis.
To create predictors derived from our full-curve analysis, we combined scores for $16$ distinct modes of amplitude variation from three types of curve data objects (vGRFs, apGRFs, and mlGRFs) and phase variation. Each mode is represented by either a set of amplitude PC scores or PNS phase scores. Below are listed the $16$ sets of scores that together formed our predictor variables:
@@
We developed a set of predictors based on conventional discrete summaries of GRF curves found in the literature, including the first peak of the vGRF curve (maximum over $0-50\%$ of the stance), the minimum value of the apGRF curve (minimum over $0-100\%$ of the stance) and the maximum value of the apGRF curve (maximum over $0-100\%$ of stance).
Using JSW as the dependent variable, we conducted a comparison of multiple linear regression models. We first used either our full-curve predictors or the set of conventional discrete summaries as predictor variables. Since we treat curves as data objects, we used all 2686 gait curves as individual observations, where each curve was represented either by its full-curve predictors or by its corresponding set of conventional discrete summary predictors. We adjusted our hypothesis testing inference to account for the dependence among curves from the same patient, as described in the next two paragraphs.
The model using full-curve predictors achieved a coefficient of determination $R^{2}=0.2$, whereas the model using conventional discrete summary predictors had a much smaller value of $R^{2}=0.024$. Although both $R^{2}$ values were small, the model using full-curve predictors demonstrated a substantially stronger association with JSW. Given that OA is a complex, multi-system disease influenced by metabolic, inflammatory, post-injury, and aging-related factors beyond biomechanics, modest $R^{2}$ values are expected and common in the field.
We assessed the significance of the $R^{2}$ values using a bootstrap-based F-test. Standard F-tests assume independent observations, which is violated in our dataset due to a clear dependence between the multiple gait curves per patient. To account for this, we implemented a bootstrap procedure in which patients were resampled with replacement, and all gait curves associated with a selected patient were included in the resampled dataset. For each of 1000 bootstrap resamples, we refit the models and computed the F-statistics, constructing null distributions for model fit. All other bootstrap analyses in this paper also used 1000 replications. The significance of each model’s $R^{2}$ value was then assessed by computing the proportion of bootstrap F-statistics greater than or equal to the observed F-statistic from the original data. This resampling approach reflects the natural dependence structure in the data, providing an appropriate null distribution for inference. The results of this bootstrapping procedure showed that the $R^{2}$ for the model with full-curve predictors was significantly different from zero (p-value $0.022$), while the $R^{2}$ for the model with conventional discrete summary predictors was not significantly different from zero (p-value $0.22$).
Considering the complexity of OA disease and that the $R^{2}$ measure may arbitrarily favor larger models, a more nuanced comparison is appropriate to confirm the value added by using full-curve predictors. With this motivation, we conducted F-tests for nested models via bootstrapping to provide a more robust comparison, detailed in the following section.
@@ 4.1 Nested model comparison
We assessed the value added by full-curve predictors over conventional discrete summaries by comparing nested linear regression models. Specifically, we evaluated whether full-curve predictors provided additional predictive value beyond conventional discrete summaries using a bootstrap-based F-test approach.
@@
To compare these models, we computed the F-ratio for two cases: (1) a full model containing both full-curve predictors and conventional discrete summaries versus a reduced model containing only the full-curve predictors, and (2) the same full model versus a reduced model containing only the conventional discrete summaries. The null hypothesis for this F-test states that the additional predictors in the full model have coefficients equal to zero, meaning they do not contribute additional explanatory power. Failing to reject the null hypothesis when the reduced model contains the full-curve predictors suggests that the conventional discrete summaries do not add predictive value beyond the full-curve predictors. Conversely, rejecting the null hypothesis when the reduced model contains the conventional discrete summaries indicates that the full-curve predictors offer additional predictive value not captured by conventional discrete summaries alone. Similar interpretations apply when reversing the roles of the full-curve and conventional discrete summaries.
As before, we used all 2686 gait curves as individual observations, representing each curve either by its full-curve predictors or its corresponding set of conventional discrete summary predictors. We again accounted for the dependence between observations using a patient-level bootstrap resampling procedure. We then fit the full and reduced models to each resampled dataset and computed the F-statistic, constructing a null distribution for model comparison.
The results showed that the reduced model with only full-curve predictors was not rejected at the $0.05$ significance level (p-value $0.34$), suggesting that the addition of conventional discrete summaries did not improve model performance, whereas the reduced model with only conventional discrete summary predictors was rejected (p-value $0.016$), indicating that full-curve predictors provided additional predictive value beyond conventional discrete summaries. This comparison highlights that full-curve predictors offer greater predictive performance and explain more variation in the important disease indicator, JSW.
As previously noted, JSW is just one measure of OA severity. In the next section, we examine the other key indicators - KLG and WOMAC pain and function scores - along with additional clinical measures relevant to OA. By analyzing these broader severity measures and clinical traits, we further assess the added value of full-curve predictors compared to conventional discrete summaries.
4.2 Predicting clinical traits
Various clinical traits relevant to patient health, OA severity, and pain/functionality were collected as part of the IDEA study. These included data from the participant level on sociodemographics, pain, mental and physical health, OA severity, and others. We used a subset of these clinical traits as dependent variables to compare full-curve predictors with conventional discrete summaries using the same nested model procedure as before. We selected 40 out of 72 clinical traits for analysis by excluding those that were redundant with other traits and those with a high number of missing values.
Figure 15 in the supplementary material shows a heat map of missing values after removing redundant traits, with blue lines indicating missing values for each clinical trait. The first 40 traits in that figure are the selected ones. Recall that the last column in that figure corresponds to JSW, which is analyzed separately in Sections 4 and 4.1. Since additional data are available for other clinical traits, we conducted the analysis of disease severity using JSW separately, as discussed in those sections. Since these other clinical traits had relatively few missing values, we imputed the remaining missing values using the mean value of each corresponding trait. Table 1 in the supplementary material provides an overview of the selected clinical data, including abbreviated trait names, brief descriptions, and trait types.
For each additional clinical trait, we repeated our comparison of multiple linear regression models, with the trait as the dependent variable and all 2686 gait curves represented either by full-curve predictors or conventional discrete summaries as independent variables. We maintained a consistent modeling strategy across all traits without specifically adjusting for binary or ordinal response variables or potential multicollinearity. While multiple linear regression may be conservative in such cases, our goal was to assess the relative predictive value of full-curve versus conventional discrete summary predictors under a uniform framework rather than to maximize predictive power for any specific trait. When conducting inference for each trait model, we again accounted for the dependence between observations using a patient-level bootstrap resampling procedure.
Figure 6 presents a scatter plot in which each axis represents the $R^{2}$ value for models using the respective predictor sets. The $R^{2}$ values were not significantly different from zero for both models using full-curve predictors and those using conventional discrete summary predictors when predicting the following clinical traits: Positive affect, Negative affect, Life satisfaction, SF-36 Mental, Mini Mental State, Number of falls, Heart attack, Heart failure, Heart arrhythmia, Stroke, Hypertension, Cancer, Diabetes, Depression, Pacemaker, Marital status, Education, Alcohol use, and Smoking. These traits, represented by gray circles in the figure, include comorbidities, mental and physical health measures not necessarily related to OA, and sociodemographic variables unlikely to be directly linked to biomechanics.
In contrast, the $R^{2}$ values were significantly different from zero for both predictor sets when modeling traits more directly relevant to gait, including: Gait efficacy, Balance confidence, Exercise adherence, Fitness satisfaction, WOMAC function, SF-36 Physical, Hip circumference, Walking speed, Walk distance, BMI, Fear of falling, Thinking about falling, and Male. These traits are represented by green stars in the figure.
Notably, several traits had an $R^{2}$ value significantly different from zero for models using full-curve predictors but not for models using conventional discrete summary predictors. These traits include: Worst knee KLG, WOMAC pain, Weight, Waist circumference, Age, CRP, IL-6, and Black. These variables are all potentially important in knee OA. In particular, Worst knee KLG and WOMAC pain represent two important measures of OA severity, as discussed earlier, while CRP and IL-6 are inflammatory biomarkers that may contribute to OA development and progression. These traits are represented by a green plus sign in the figure. Importantly, there were no traits for which models using conventional discrete summary predictors had an $R^{2}$ value significantly different from zero while models using full-curve predictors did not.
Figure 6: Scatter plot of $R^{2}$ values for models using full-curve predictors versus conventional discrete summaries. Gray circles indicate traits where $R^{2}$ was not significantly different from zero for either predictor set. Green stars mark traits where both models had significant $R^{2}$ values, while green plus signs indicate significance only for full-curve predictors. No traits showed significant $R^{2}$ values for conventional discrete summaries alone. All points fall below the diagonal ($x=y$), demonstrating consistently higher $R^{2}$ values for full-curve models. Points farther from the diagonal reflect greater differences, with the largest gaps observed for Walking speed, Walk distance, Male, Hip circumference, BMI, Waist circumference, Weight, and Black, which are labeled in the figure.
All points lie below the diagonal line ($x=y$), shown as a black dashed line in the figure, indicating that $R^{2}$ values for models based on full-curve predictors are consistently higher than those for conventional discrete summary predictors. The distance from the diagonal reflects the magnitude of this difference, with points farther from the line representing greater improvements in predictive performance when using full-curve predictors. As labeled in the figure, the largest differences in $R^{2}$ values were observed for models predicting Walking speed, Walk distance, Male, Hip circumference, BMI, Waist circumference, Weight, and Black.
Despite generally low $R^{2}$ values for both predictor sets, full-curve predictors consistently explain more variation in clinical traits, indicating that they capture more information than conventional discrete summaries.
4.3 Nested model comparison for predicting clinical traits
A nested model comparison again provided further insight into the added value of full-curve predictors. We computed the F-ratio for a full model containing both full-curve and conventional discrete summary predictors versus reduced models containing only one predictor set. We then generated a null distribution for the F-tests via patient-level bootstrapping and performed the inference as described previously. When no bootstrap replications were larger than the original statistic, we defined the p-value to be $\frac{1}{2*1000}=0.0005.$
@@ 4.3 Nested model comparison for predicting clinical traits
Figure 7: Scatter plot of bootstrap p-values resulting from F-tests of nested models, shown on a logarithmic scale. The horizontal axis represents the p-values for testing reduced models with full-curve predictors, and the vertical axis represents those for testing reduced models with conventional discrete summary predictors. For readability, the axes are labeled on the original scale. A red dashed line marks the $\log_{10}(0.05)$ threshold on both axes, corresponding to the $0.05$ significance level. Gray circles represent traits where neither reduced model was rejected at this significance level, indicating that neither predictor set adds substantial predictive value. Green plus signs represent traits where the reduced model with full-curve predictors was not rejected, but the reduced model with conventional discrete summary predictors was rejected, highlighting the added value of full-curve predictors. No points appear to the left of the vertical red dashed line, indicating that the reduced model with full-curve predictors was never rejected and is consistently a suitable approach.

### v3

@@ 3.1.1 Implementation and penalty parameter selection
To address the lack of a common underlying structure, we registered the full dataset using the penalized elastic shape analysis procedure. By adjusting the elasticity parameter $\lambda$, we aligned the full set of GRF curves without distorting the shape of the atypical curves. To determine an appropriate $\lambda$, we computed warping functions iteratively over a grid of candidate values and visually examined the alignment results, with particular attention to atypical cases. We refined the grid in regions where $\lambda$ values produced reasonable results, selecting values that avoided excessive smoothing of features in the atypical curves while still capturing meaningful phase variation. Over-alignment led to sharp corners forming a staircase-like pattern in the warping functions, indicating drastic stretching and compression of curves with differing underlying structures, such as unimodal curves.
@@
Based on this evaluation, we selected $\lambda=2$ for mitigating the staircase effect without totally sacrificing alignment. To illustrate the alignment trade-off, Figure 3 compares results for three $\lambda$ values: the fully elastic case $\lambda=0$ in the top row, our selected $\lambda=2$ in the middle row, and a less elastic case $\lambda=4$ in the bottom row. Atypical cases are colored by speed, while other curves are shown in gray, as in Figure 2. The leftmost panels display the warping functions for each $\lambda$, where $\lambda=0$ results in sharp staircase-like patterns indicative of over-alignment, while $\lambda=4$ yields warping functions tightly clustered around the identity, suggesting insufficient alignment. The second through fourth columns show the aligned curves. Results for additional values of $\lambda$ can be found in Figure 11 of the supplementary material.
@@
While our approach treated alignment as agnostic to a specific performance criterion, a more formal selection method could be constructed depending on the analytical objective. For example, if the interest is in evaluating alignment quality in reference to a predictive or inferential task, $\lambda$ could be optimized via cross-validation to maximize a relevant performance metric. However, in finite samples, cross-validation is prone to noisy selection due to its slow convergence to optimal results (see Hall and Marron (1987) in the context of bandwidth selection for kernel density estimation). Alternatively, if the goal is to estimate an underlying common signal, Kim et al. (2023) provides a scale-space approach for estimating both the shape of the unknown signal and the signal itself. Defining a notion of optimality depends on the specific context of the analysis, and the choice of $\lambda$ for different analytical objectives is an open question for further research.
@@
Figure 3: Alignment results for three elasticity parameter values: fully elastic ($\lambda=0$), the selected $\lambda=2$, and a more rigid case ($\lambda=4$). Atypical cases are colored by speed, while other curves are shown in gray. The leftmost panels display the warping functions, and the second through fourth columns show the aligned curves. The staircasing effect in the fully elastic case ($\lambda=0$) indicates over-alignment, while $\lambda=4$ produces warping functions tightly close to identity, suggesting insufficient alignment. The selected $\lambda=2$ balances these effects.
@@
Next, we obtained amplitude objects by applying the (common) set of warping functions to each set of original GRF curves, which provides an intuitive representation of amplitude in each component. The amplitude objects obtained for different values of $\lambda$ are shown in the second through fourth columns of Figure 3.
@@ 3.2 Modes of variation
In OODA terminology, a collection of members of the object space that summarize one component of variation and is in some sense one-dimensional is called a mode of variation. For example, in the vector matrix case, a mode of variation is a rank-one matrix. We can obtain modes of variation through Principal Component Analysis (PCA), where each object is considered as a point in high dimensional space (column vector). For an introduction to PCA, see Jolliffe (2002). Amplitude modes of variation (Tucker et al., 2013) were obtained for each direction of the ground reaction force (vGRF, apGRF, mlGRF) computing PCA on the set of 2,686 Fisher-Rao aligned curves, each corresponding to an individual gait observation. The sets of input curves are shown in the second, third, and fourth columns of the middle row of Figure 3.
@@
Figure 4 shows the modes of variation of the amplitude objects of the vGRF, where the curves are colored according to walking speed. The first mode of variation, shown in the first panel of the second row of that figure, is associated with walking speed and reflects the contrast in peak heights and valley depths. Faster walkers (indicated in red in the rainbow color scheme) generally exhibit higher peaks and lower valleys, while slower walkers (purple in the rainbow color scheme) have lower peaks and a shallower valley. The middle column of that figure displays the largest (dashed curve) and smallest (dotted curve) PC projections added back to the mean curve, which is shown as a solid black curve in each of the middle panels. The middle panel of the first mode shows that the slowest walkers (dotted curve) exhibit a vertical amplitude force that appears unimodal and does not exceed body weight (1 on the vertical axis), indicating that these walkers do not fully transfer their weight to the striking limb. This type of gait can be thought of as “shuffling.” The second and third modes of variation are about the second and first peak, respectively. The middle panel in the third row shows the largest and smallest PC2 projections added back to the mean curve, distinguished with a dashed (largest) and dotted (smallest) line type. These extremes show that variation in this mode is mostly in the height of the second peak. In the panel below, the extremes of the third mode indicate phase variation in the first peak that is unique to the vertical component. The fourth mode of variation reflects the overall magnitude, particularly in the mid-stance phase. The second, third, and fourth PC projection extreme curves all suggest that some curves have a small third bump before the first peak. In gait analysis, this pattern is known as the heel-strike transient (HST), a rapid and transient rise in the vGRF immediately after ground contact. As discussed in Blackburn et al. (2016) and references therein, the presence and characterization (e.g. magnitude) of HST can indicate impulsive loading, which influences cartilage degradation and symptoms of OA. However, Blackburn et al. (2016) also noted that methods for identifying HST can be unreliable. The amplitude modes of variation we identified offer a potentially viable method for reliably identifying and understanding the HST.
@@
We applied the PNS-based functional PCA methodology to the set of (common) warping functions to obtain phase modes of variation. We found that the great sphere decomposition from PNS yielded the most interpretable phase modes of variation because of weak interpretability of small sphere variation. Figure 5 depicts an intuitively useful notion of phase variation represented by warpings of the Karcher mean of the vGRF curves. The warping functions used to create these visualizations were generated by taking the inverse of the phase PNS projections added to the 45-degree line (identity warp). In each panel of the figure, the curves are colored based on the PNS scores for the corresponding mode, with cyan indicating the lowest scores and magenta indicating the highest. It is important to note that the curves are plotted in the order of the corresponding score, as over-plotting is an issue. The first mode (first panel) shows an overall shift in timing, with most apparent differences in the timing of the first peak (maximum heel-strike force) and valley. The second mode in the next panel appears to explain variability in the closeness of the peaks: the cyan curves are the curves with peaks closer together and the magenta curves have peaks farther apart. The third mode represents an overall phase shift (left vs. right) and seems to suggest that curves having a small third bump before the first peak correspond with earlier timing (cyan curves), especially an earlier second peak. The fourth mode appears to highlight variability in the timing of the second peak, independent of the rest of the curve.
@@ 4 Comparison to Conventional Discrete Summaries
We investigated the added value of analyzing patterns across the entire movement curve, rather than relying on conventional discrete summaries of GRF curves. For this purpose, we compared how strongly full-curve modes versus conventional discrete summaries were associated with OA-related clinical traits using a nested regression framework, detailed in Section 4.2. While some of the traits we considered are clinically meaningful to predict from gait biomechanics, the aim of this analysis was to quantify and compare the strength of associations between OA-related traits and gait features derived from full-curve analysis versus conventional discrete summaries. In this section, we outline the sets of gait features that served as independent variables in our models, and in the following section, we introduce the OA-related clinical traits.
To create independent variables derived from our full-curve analysis, we combined scores for $16$ distinct modes of amplitude variation from three types of curve data objects (vGRFs, apGRFs, and mlGRFs) and phase variation. Each mode is represented by either a set of amplitude PC scores or PNS phase scores. Below are listed the $16$ sets of scores that together formed our full-curve independent variables:
@@
We developed a set of independent variables based on conventional discrete summaries of GRF curves found in the literature, including the first peak of the vGRF curve (maximum over $0-50\%$ of the stance), the minimum value of the apGRF curve (minimum over $0-100\%$ of the stance) and the maximum value of the apGRF curve (maximum over $0-100\%$ of stance).
4.1 OA Clinical Traits
The IDEA study defined several OA disease outcomes and symptoms of interest (Messier et al., 2009, 2013). These included mechanistic outcomes: knee joint compressive force and inflammatory biomarkers (interleukin-6 [IL-6] and C-reactive protein [CRP]); and clinical outcomes: self-reported pain and function, mobility, and health-related quality of life. Increased knee joint compressive force is known to contribute to cartilage stress and degeneration and has been associated with patterns in gait biomechanics (D’Lima et al., 2012). Elevated IL-6 and CRP levels are linked to chronic inflammation and have been associated with knee OA (Messier et al., 2009). Pain and function were measured using the Western Ontario and McMaster Universities Arthritis Index (WOMAC) (Alexandersen et al., 2014). Mobility was assessed using walking speed and distance walked in a 6-minute trial, while health-related quality of life was evaluated using the SF-36 Physical and Mental Component Scales. These IDEA study outcomes were included as dependent variables in our regression analyses.
Although not specified as outcomes in the IDEA study, we analyzed additional biomedical measures that have been examined in other OA research. Notably, radiographic OA severity is evaluated using joint space width (JSW) and Kellgren–Lawrence grade (KLG), and both are key metrics for diagnosing and monitoring disease progression. Prior studies have reported associations between gait features and variability in these structural measures (e.g. Kwon et al. (2019); Jansen et al. (2024)), making them particularly relevant to this analysis. OA is also associated with elevated fall risk, and existing studies have examined fall risk as an outcome in patients with knee OA (for example, Rosadi et al. (2022)). Furthermore, anthropometric characteristics (e.g., body weight) and sociodemographic traits (e.g., age, sex, race) are important factors in OA risk, symptoms, and treatment (see, for example, Sims et al. (2009); Allen et al. (2022); Chang et al. (2024)). These additional measures, together with the IDEA study outcomes, are summarized in Table 1 of the supplementary material and are collectively referred to as clinical traits throughout the paper. For limb-level measurements (such as joint compressive force, KLG, and JSW), we used the patient-level clinical aggregates of these measures reported in the IDEA study for our analysis (Messier et al., 2009, 2013).
Since the IDEA study was an 18-month clinical trial of interventions in patients with advanced knee OA, traits were collected at different intervals throughout the study. For the purpose of this analysis, focusing on baseline values provides the most consistent basis for comparison and may better reflect OA variability in the dataset. For example, although change in JSW is a key marker of progression, prior studies suggest that knees at similar OA stages progress at comparable rates (Benichou et al., 2010), making baseline JSW a more meaningful indicator of disease severity for the IDEA participants. Additionally, using baseline values avoids the need to account for treatment effects, which are not of interest in this analysis.
Figure 15 in the supplementary material shows a heat map of missing values in the baseline data, with blue lines indicating missing entries for each trait; note that JSW has a relatively high rate of missing data. Baseline JSW measurements were missing for 126 of the 454 IDEA participants. Although the reasons for these missing values were not reported in the study, comparison of the other clinical traits between participants with and without baseline JSW showed no apparent sampling bias. This left 328 participants with JSW data for our analysis. Due to the large proportion of missing values, analysis involving JSW was restricted to the complete subset, with gait variables subset accordingly. Missing values for other traits, which were relatively few, were imputed using the mean of each trait (or the mean rounded to the nearest integer for integer-valued data). Importantly, there were no missing values for traits that may be particularly sensitive to imputation, such as categorical traits like sex and race, and low-resolution ordinal measures like KLG. The treatment of different types of dependent variables (count, binary, ordinal, continuous) for modeling purposes is discussed in detail in the following section.
@@ 4.2 Nested Model Comparison
We assessed the added value of full-curve gait analysis over conventional discrete summaries by comparing nested regression models. Specifically, we evaluated whether full-curve models provided additional explanatory value beyond conventional discrete summaries using a bootstrap-based likelihood ratio test (LRT) approach. Since we treat curves as data objects, all 2,686 gait curves were used as individual observations, except when modeling JSW, which was analyzed on the subset described earlier. Each trait was treated as a dependent variable and a full model fit on the combined set of gait features (full-curve modes plus conventional discrete summaries), while reduced models were fit using either the full-curve modes alone or the conventional discrete summaries alone.
The choice of regression model depended on the type of dependent variable: count, binary, ordinal, or continuous. Number of falls was the only count-valued trait and was modeled using Poisson regression. Sex and race were the only binary-valued traits and were fit using logistic regression. Ordinal logistic regression was used for fall-related traits, WOMAC pain and function, health-related quality of life measures, and KLG. KLG is a radiographic score of OA severity, while the other ordinal traits represent scores from Likert-type scales. All remaining traits were continuous and fit using linear regression. As detailed in the previous section, missing trait values (except for JSW) were imputed using the trait mean. For integer-valued traits, the mean was rounded to the nearest integer, whereas for continuous traits it was used directly.
@@
To compare nested models, we computed likelihood ratio tests for two cases: (1) a full model containing both full-curve modes and conventional discrete summaries versus a reduced model with only the full-curve modes, and (2) the same full model versus a reduced model with only the conventional discrete summaries. The null hypothesis in each case is that the reduced model explains the data as well as the full model, while the alternative is that the additional independent variables in the full model significantly improve model fit. Failing to reject the null when the reduced model contains the full-curve modes suggests that discrete summaries add no explanatory value beyond the full-curve modes. Conversely, rejecting the null when the reduced model contains the discrete summaries is evidence that full-curve modes provide additional explanatory value not captured by discrete summaries. The same logic applies when the roles of the two variable sets are reversed. Note that for linear regression models, the likelihood ratio test is equivalent to the nested-model F-test.
Typically, a LRT statistic is compared to a chi-squared reference distribution, but this relies on an asymptotic result which assumes independent observations. This assumption is violated in our dataset due to a clear dependence between the multiple gait curves per patient. To account for this, we implemented a bootstrap procedure in which patients were resampled with replacement, and all gait curves associated with a selected patient were included in the resampled dataset. For each of 1,000 bootstrap resamples, the regression models were on the selected sample and the corresponding LRT statistic was computed. This resampling approach reflects the natural dependence structure in the data and generates an appropriate reference distribution for inference. All other bootstrap analyses in this paper also used 1,000 replications. The significance of the observed LRT statistic from the original data was then assessed by computing the proportion of bootstrap LRT statistics greater than or equal to the observed LRT statistic. When no bootstrap replications were larger than the original statistic, we defined the p-value to be $\frac{1}{2*1000}=0.0005.$
@@
Figure 6: Scatter plot of bootstrap p-values resulting from LRTs of nested models, shown on a logarithmic scale. The horizontal axis represents the p-values for testing full-curve reduced models, and the vertical axis represents those for testing conventional discrete summary reduced models. For readability, the axes are labeled on the original scale. Gray circles represent traits where neither reduced model was rejected at this significance level, indicating that neither variable set adds substantial explanatory value beyond the other. Green plus signs represent traits where the full-curve reduced model was not rejected, but the conventional discrete summary reduced model was rejected, highlighting the added value of the full-curve approach. Note that the full-curve reduced model was never rejected and is consistently a suitable approach.